OpenStudy (anonymous):

Solve for the variable in this equation. x/2+1/4=3/4

OpenStudy (anonymous):

First subtract 1/4 from both sides. Than multiply both sides by 2 and you got your variable.

OpenStudy (anonymous):

I dont know how to subtract from both sides

OpenStudy (anonymous):

what's 1/4 - 1/4

OpenStudy (anonymous):

0

OpenStudy (anonymous):

great you did the left side now lets do the right side whats 3/4 - 1/4

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

hmm when adding or subtracting fraction you need to add or subtract the top number. what's 3 - 1 ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

great the top number is 2 and the bottom stays the same. 2/4 Thats what you get. ok now you have x/2 = 2/4 (remember the left side you canceled the 1/4 ) now multiply both sides.

OpenStudy (anonymous):

whats x/2 times 2

OpenStudy (anonymous):

4

OpenStudy (anonymous):

well no it's actually (2x)/2 ... the 2s in both the top and bottom will cancel so you are left with just x now on the right side you do the same 2/4 times 2 = (2*2)/4 Whats 2 times 2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ok whats 4/4

OpenStudy (anonymous):

16

OpenStudy (anonymous):

umm not your multiplied it but we want to divide. try again 4 divided by 4?

OpenStudy (anonymous):

oh 1

OpenStudy (anonymous):

great so the right side is 1

OpenStudy (anonymous):

and the left side is x

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Oh thx

OpenStudy (anonymous):

Sorry to interrupt you guys, How old are you PaigeMya?

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

Age doesn't matter....

OpenStudy (anonymous):

That's a creepy question @amogh

OpenStudy (anonymous):

Yeah why do you want to know my age?

OpenStudy (anonymous):

Oh don't take me wrong. Just wanted to know because everyone needs different treatment , clarification needed. And yeah Age doesn't matter, you're right, but that's the best I could ask concerning what I wanted.

OpenStudy (anonymous):

Okay well im ten

OpenStudy (anonymous):

Can i also just ask another question right now for you guys to help me with?

OpenStudy (anonymous):

It's not about the age but the level of math you are taking.

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

@Romero Exactly, but how can someone answer to that?

OpenStudy (anonymous):

Thanks, find the lowest common denominator that would eliminate the fractions in this equation. 8/9x+2/3x=1/9x+45

OpenStudy (anonymous):

By the type of question they asked. Someone can ask a linear algebra question and they can be a freshman in college.

OpenStudy (anonymous):

Which fraction has the lowest top number (numerator)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1/9

OpenStudy (anonymous):